A party rental company has chairs and tables for rent. The total cost to rent
4
chairs and
8
tables is
$57
. The total cost to rent
2
chairs and
3
tables is
$22
. What is the cost to rent each chair and each table?

Answer by stanbon(75887)

A party rental company has chairs and tables for rent. The total cost to rent
4 chairs and 8 tables is $57
--------
The total cost to rent 2 chairs and 3 tables is $22.
--------------
What is the cost to rent each chair and each table?
Equations::
4c + 8t = 57
2c + 3t = 22
----------------
Modify for elimination:
4c + 8t = 57
4c + 6t = 44
---------------------
Subtract and solve for "t":
2t = 13
t = $6.50 (cost of one table)
-----
Solve for "c":
2c + 3t = 22
2c + 19.5 = 22
2c = 2.5
c = $1.25 (cost of one chair)
--------------
Cheers,
Stan H.
------------